A satellite of mass 200kg is placed into Earth orbit at a height of 200km above the surface.
(a) Assuming a circular orbit, how long does the satellite take to complete one orbit?
(b) What's the satellite's speed?
1
Expert's answer
2021-02-10T10:11:12-0500

Answer

b) r=Re+h

Me massof earth

In circular motion

Centripetal force is equivalent to gravitational force

"F_c=F_g"

"\frac{mv^2}{r}=\frac{GmM_e}{r^2}"

"v=\sqrt{\frac{GM_e}{r}}"

Putting all value according to the question

"v=7.79\times10^3" m/s

a) now time taken to complete one orbital motion

"T=\frac{2\pi r}{v}"

Putting value of r and v

T=1.48hrs.